OpenStudy (anonymous):

Solve the equation. If your answer contains a fraction, use '/' to represent the fraction bar. If your answer is a mixed number, put a space between the whole number and the fraction. For example, write 5 1/2 as 5 1.2 3/7b = 22 b=?

OpenStudy (anonymous):

3/7b = 22 b = 22 * 7/3 = 154/3